If a sample of gas occupies 6.8 liters at 327 degrees celcius what is the volume at 27 degrees celcius if the pressure does not change?
1
Expert's answer
2013-03-15T09:50:44-0400
This question can be solved using the ideal gas law:

PV = nRT
P is pressure
V is volume
n is amount
R is gas constant = 8.31
T is temperature

If V and T are variable, P,n,R are constants

P1V1 = nRT1
P2V2 = nRT2

V1/T1 = nR/P1
V2/T2 = nR/P2
P1 = P2

V1/T1 = V2/T2

Now it is possible to find V2

V2 = V1*T2/T1 = 6.8*(273+27)/(273+327) = 3.4 L